OpenStudy (anonymous):

Let n = 3^17 + 3^10 It is known that 11 divides into n+1. If n can be written in base 10 as ABCACCBAB, where A,B,C are distinct digits such that A and C are odd and B is not divisible by 3, find 100A + 10B + C.

OpenStudy (mathmath333):

first find \(A\) and \(B\) from \(\large \color{black}{\begin{align} n\pmod {100}\hspace{.33em}\\~\\ \end{align}}\) ABCACCBAB then use the divisiblity criteria of \(11\) that \(\large \color{black}{\begin{align} B+A+C+A-(A+C+C+B+(B+1))=11n\hspace{.33em}\\~\\ \end{align}}\) where \(n=0,1,2,\cdots ,+\infty\)

OpenStudy (mathmath333):

does this makes sense

OpenStudy (anonymous):

:-) not much Sir, trying out other things

OpenStudy (mathmath333):

difficulty in \(n \pmod {100}\) ?

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

\(\large \color{black}{ \dfrac{3^{17} + 3^{10}}{100}\hspace{.33em}\\~\\ \implies \dfrac{3^{10}\left(3^{17} + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{9^{5}\left(3^{7} + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{(10-1)^{5}\left(3^{7} + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{\binom{5}{0}\cdot (-1)^5\cdot(10)^0+\binom{5}{1}\cdot (-1)^4\cdot(10)^1+\cdots )\left(3^{7} + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{(-1+50 )\left(3^{7} + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{(49 )\left(3^{7} + 1\right)}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{3^{7}}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{3\cdot 3^{6}}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{3\cdot 9^{3}}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{3\cdot 729}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{3\cdot 29}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{87}{100}\hspace{.33em}\\~\\ \color{red}{\implies}\dfrac{-13}{100}\hspace{.33em}\\~\\ \implies \dfrac{(49 )\left(-13 + 1\right)}{100}\hspace{.33em}\\~\\ \implies \dfrac{-588}{100}\hspace{.33em}\\~\\ \implies \dfrac{-88}{100}\hspace{.33em}\\~\\ \implies \dfrac{12}{100}\hspace{.33em}\\~\\ \LARGE \implies 12\hspace{.33em}\\~\\ A=1,B=2\hspace{.33em}\\~\\ }\)

OpenStudy (mathmath333):

\(\large \color{black}{ B+A+C+A-(A+C+C+B+(B+1))=11n\hspace{.33em}\\~\\ 2A+B+C-(A+2C+2B+1)=11n\hspace{.33em}\\~\\ A-B-C=11n+1\hspace{.33em}\\~\\ 1-2-C=11n+1\hspace{.33em}\\~\\ C=-11n-2\hspace{.33em}\\~\\ \normalsize \text{u want C as single digit positive number so}\hspace{.33em}\\~\\ C=-11(-1)-2\hspace{.33em}\\~\\ C=9\hspace{.33em}\\~\\ }\)

OpenStudy (anonymous):

A=1,B=2 and C=9.